Inserting parentheses with
My problem
60-18÷2+4+10÷2=12

Found 2 solutions by ikleyn, Edwin McCravy:
Answer by ikleyn(52776)

.

    (60-18) divided by (2+4) + (10 divided by 2) = 12.

Answered, solved and explained.

And completed.


---------------
comment from student: Thx What is rule of thumb inserting parentheses Another 2 problems need to help my 10 yeAr old thx ur great

80-36 / 2 + 6 - 2 x 5 + 1 =44 3 x 5 + 4 - 2 x 2 + 6 - 4 = 15
---------------


My response.

First of all, there is NO any rule of thumb and/or any other rule, at all . . .

The method is "trial and error": you TRY, then you LOOK, then you TRY again . . . and so on . . .


Second thing is that this window for your comments, through which you posted your comment, IS NOT DESIGNED
for posting new problems to some (to any) tutor in this forum.

This window is designed and intended for different purposes . . .

    I hesitate to say: it is intended mainly for visitors to say "Thanks . . .  "

So, you better re-submit your last request to the forum by a regular way, as you do it usually.

Other tutors also want to participate, and I can not deprive them this joy.


The third thing is that such problems are ENTERTAINMENT problems.

Their destination is to provide a joy of solution to people of all ages, from 7 to 77 (or from 8 to 88).
Therefore, give this possibility to your child to play with parentheses - he will be grateful to you . . .



Answer by Edwin McCravy(20054)

(60-18)÷(2+4)+10÷2 = 12
------- ----- ----
   |   |  |  |  |  
   42  ÷  6  +  5
   --------
       |     |  
       7     +  5
       ----------
             | 
            12

Edwin